# Hey Dashboard!: Supporting Voice, Text, and Pointing Modalities in Dashboard Onboarding

### v1

@@ 6.2.3 Hard Tasks
Hard exploratory and interpretive tasks were the most demanding, often involving multiple steps and rephrasing. Participants engaged in multi-modality sequences (e.g., mouse → voice → keyboard). P2, P3, P4, and P6 used the mouse and voice modality and rephrased the questions in their own way to interact with Diana. P2, P3, and P6 asked several follow-up questions using the voice modality, indicating that as tasks became more complex, the need for human-like onboarding became more prevelant. P1 and P5 switched between keyboard and other modalities. Despite the use of all the modalities, P1 could not solve one of the hard tasks, indicating that the instruction prompt to Diana could have been improved. This was the only unsolvable task during the entire user study session across all participants. P2 also used the mouse modality, utilizing the radial menu feature that helped him confirm his answer to one of the tasks.
6.2.4 Preferred Modalities and Features
As the participants performed the user study, their choice of preferred modality became clear. Despite varying levels of prior experience, all participants experimented with speech interaction. Even those with no prior exposure (P1, P2, P6) adopted it during the study, with P2 initially hesitant but later describing it as “comfortable, almost not robotish.” Voice was valued for its speed and natural phrasing, while chat was preferred for step-by-step guidance (P1 and P5). Mouse interactions, such as lasso selection, were consistently used to define a focus before turning to another modality (P3 and P4). Across modalities, the visual highlight feature was universally appreciated to orient user attention.
6.2.5 Underused Modalities and Features
The radial menu was the least used feature. While P2 found it helpful for confirmation, most others described it as unintuitive, overly detailed, or unfamiliar: “I didn’t know what specifically to use it for […] the surface things were good enough” (P4). The user guide was largely ignored, with P1 and P3 admitting that they forgot it existed. P6 said she would have preferred to read it if time allowed. This indicates that static documentation can be easily overshadowed by interactive support.
6.3 Reflection on the System
@@
In the post-task interview, participants reflected on three main themes: their choice of modality, their experience with the onboarding, and the comparison to previous onboarding experiences.
6.3.1 The Value of Multimodality.
@@
Several participants emphasized that having multiple options was useful, as it allowed them to adapt depending on time, task, or personal style. P1 summarized this flexibility: “If you are in a hurry, you can just talk to it and it goes faster. If you have time to sit down and think, then you can use the chat and the radial menu… But I would probably use writing at most, because I’m old-fashioned.”
Similarly, P2 appreciated having “fallbacks” between modalities. P3 and P4 also found the use of multiple modalities, especially mouse and voice quite helpful, with P4 saying that she “definitely liked the combination of selecting [lasso] and speech.”
6.3.2 Strengths of Voice-based Interaction and Visual Highlights Feature.
@@
Speech interaction was well received, even among those without prior experience. P2 remarked that “it was very comfortable voice, I could hear that it was mechanical, but it was not that mechanical.” He also appreciated the real-time response, and when paired with the visual highlighting feedback, it helped him see what Diana was saying in response to his question (DG4). As P2 was an expert, he could verify the correctness of the answer provided by Diana. As the answer was phrased naturally and accurately, he said “and when it did that [gave him the right steps to the answer] it had my trust, that I could actually ask the voice interaction a question.” P3 described the voice modality and the response from Diana informative and that he “really [had] fun interacting with her,” while P4 appreciated that speech “doesn’t stop when you are doing something else on the dashboard.” The mouse modality, in the form of lasso selection, was also valued when combined with voice, as the combination helped participants anchor their queries to specific visuals.
6.3.3 Limitations of Chat and Radial Menu.
Although useful for step-by-step guidance, the chat via keyboard was sometimes described as slow or effortful (“chatting took too long to write compared to just speaking,” P2). The radial menu was the least favored feature: while P2 found it helpful for confirming a difficult task, others found it overwhelming or unfamiliar. P4 explained, “It had more deep details that I normally wouldn’t look for… Unfamiliarity with it made me use it less.”
6.3.4 Comparison to Prior Onboarding.
@@
Compared to step-by-step onboarding methods, Diana was described as more flexible and empowering. P4 compared it to video game tutorials where you are forced to proceed linearly: “You have to explore one step at a time… Here you have the freedom to onboard yourself the way you want to.” P1 echoed this, preferring to ask for help when needed rather than ”click all over” through a fixed tutorial (DG5). For P2, who regularly onboards colleagues via videos and one-on-one sessions, Diana represented a potential time saver. Although he values personal interaction, he saw the system as a ”virtual project finance controller” (DG3) that could free up some of his workload.
7 Discussion
@@
7.1 Speech Interaction as the Human Onboarder Assistant
In the main study, participants frequently preferred speech over mouse or keyboard-based interactions for onboarding help and follow-up questions. This result matches the findings of León et al. [28] on modality preferences for exploration tasks. This is also in line with the findings from Zhan et al. [47] who studied trust in healthcare voice assistants and suggested that perceived usefulness of a system is a significant factor that could explain the trust a user has in the system.
Rather than replacing human onboarders, Diana assists and augments their capabilities by handling routine questions such as how do I…? that are time-consuming to document and maintain. In the context of onboarding, the result also aligns with one of our initial goals of providing human-like support (DG3). The findings suggest that participants appreciated real-time responses and being less dependent on asking a human expert for help.
In our formative study, we also came across a surprising result: the Realtime API [2] detected that the participant was not a native English speaker and switched to another language that the person understood. While it was an unexpected behavior, the person was positively surprised about the detection. Furthermore, other participants wished for natural language interaction in their native language. While this is an option with the current large language models, it may go beyond the capabilities of a human onboarder in most cases.
7.2 Multimodal Interaction for Onboarding
@@
The less use of the contextual menu could also suggest that sophisticated visual encodings for help are not necessary for onboarding, as they were perceived as advanced or quite detailed. But, they would be extremely helpful at later stages after users have been onboarded, to allow them to access more fine-grained information, as echoed by P1.
7.3 Using Large Language Models for Onboarding
@@
While the natural interaction via voice and text was successful, it required a substantial effort at the technical level to minimize the number of hallucinations. The information sent to the LLMs was carefully designed based on the available REST APIs information and heuristics that ensure that information from the REST APIs is not wildly extrapolated. The prompts were fine tuned several times during the formative study process. We were able to significantly reduce the hallucinations related to both chat and voice based interactiions, but participants still encountered issues when they highlighted a specific section of the dashboard and then asked for guidance about it via natural language.
@@
Additionally, as the Diana has no access to the data, it could be used in any domain with ease because it only has information about the high level aspects of the visualization. It can infer trends and drivers, however, it cannot give detailed data-level information that is usually critical in business-context.
7.4 Limitations
While we included a user guide in our study to compare with standard onboarding materials, we did not compare our approach with other onboarding methods, such as in-person sessions. Future work should compare our LLM-powered approach with in-person onboarding to identify the effects on performance and user experience.
8 Conclusion
@@
We designed and developed Diana, a multimodal dashboard assistant that combines voice, text, and in-situ visual highlight and contextual help, to help users in dashboard onboarding. Through a qualitative user study, we found that participants leveraged all interaction modalities and preferred voice-based interaction over keyboard-based and mouse-based interaction. Our findings suggest that LLMs are useful for onboarding tasks when they are paired with visual affordances. It also encouraged users to ask for help without it getting into the way of their autonomy. We suggest that onboarding systems should be multimodal by design and offer a low-friction path to help users onboard themselves on dashboards.
Acknowledgments

### v2

@@ 6.2.3. Hard Tasks
Hard exploratory and interpretive tasks were the most demanding, often involving multiple steps and rephrasing. Participants many times engaged in multi-modality sequences (e.g., mouse → voice → keyboard). D2, D3, D4, and D6 used the mouse and voice modality and rephrased the questions in their own way to interact with Diana. D2, D3, and D6 asked several follow-up questions using the voice modality, indicating that as tasks became more complex, the need for human-like onboarding became more prevalent (DG2 (multimodal) and DG3 (human-like)). For instance, in Task 6, the question included the term EMEA, which is represented in the map visual on the dashboard. D3 first used the chat interface to ask about the EMEA region. However, this did not yield a meaningful explanation and only provided onboarding steps for selecting the region. For this reason, he switched to the voice interface, which provided a clear definition of EMEA.
For the same task, D2 used the voice interface and then switched to the radial interface to verify the information obtained. D5 and D6 switched between keyboard and other modalities. Despite using all modalities, D1 could not solve Task 3, which required interacting with the map and the funnel visual on the dashboard. She first rephrased the question in the chat interface, which returned a response indicating that the data was unavailable but provided onboarding instructions. However, the instructions did not explain how to select multiple points on the map. She encountered similar issues when she switched from the radial menu to the voice interface. This indicated that the instruction prompt to Diana could have been improved. This was the only unsolvable task during the entire user study session across all participants in the Diana condition.
6.3. Static Dashboard Guide (Baseline) Condition
All the participants (B1-B6) went through the provided user guide before starting the tasks. Based on their personal reading preference, a few participants (B1, B3, B5, and B6) simply glanced at the user guide, while others read the descriptions in detail (B2 and B4). B4 went through the document thoroughly and even tried out various interactions listed in the guide in orderthoroughly reviewed the document and even tried out various interactions listed in the guide to familiarize herself with the interaction concepts.
Figure 5. Participant results from the baseline condition. The results are shown for each participant (B1 - B6) broken down into task difficulty and colored by task accuracy and completeness.
The figure contains six bar charts labeled B1 through B6, each showing how many Easy, Medium, and Hard tasks were solved correctly, solved incorrectly, or not solved. Across all six charts, the overall pattern is that Easy tasks are most often solved correctly, Medium tasks show mixed outcomes with noticeable variation between participants, and Hard tasks consistently produce the greatest number of incorrect or unsolved attempts. The charts collectively illustrate a clear decline in accuracy as task difficulty increases, with each participant group exhibiting its own variation around this general trend.
6.3.1. Easy Tasks
Almost all the participants solved the easy tasks correctly. However, four out of the six participants faced difficulties due to a combination of limited dashboard familiarity, the lack of comprehensive onboarding material, and not always reading or fully processing the provided instructions. This led them to adopt workarounds. For example, B5 and B6 performed manual calculations for Task 4 because they were unsure how to find the relevant summed values directly in the dashboard. In the process, B6 discovered the bar chart tooltip unintentionally and then used it to answer the question correctly.
For Task 9, participants B3, B4, B5, and B6 manually calculated each data point, either mentally or using the provided pen and paper. Despite these efforts, B5 still gave an incorrect answer (as shown in Figure 5) because she did not realize that the tooltip displayed both the actual and the highlighted values for each point. This indicates that even basic dashboard interactions were not always transparent.
6.3.2. Medium Tasks
Medium-level tasks yielded more incorrect answers, and even participants who answered correctly often expressed uncertainty and sought verbal confirmation.
Task 1 resulted in the highest number of incorrect responses among all tasks (Figure 6). Participants B1, B2, B5, and B6 interacted extensively with the bar chart and other visuals in an attempt to locate the answer, even though the task only required a straightforward lookup and interpretation with no interaction. B3 and B4 arrived at the correct answer but were unsure about its accuracy. Task 2 also posed difficulties for B4 and B5, who provided an incorrect answer after doing a series of manual calculations. Task 7, which required selecting a time period, was particularly demanding for B2. She overlooked the time slicer entirely and focused only on the line chart showing monthly trends. She then opened the visual in single-view mode using the Power BI built-in feature and attempted to apply dataset-level filters manually. After struggling to inspect individual points in the line chart and then flipping to a table view, she eventually gave up. Other participants solved the task, though B6 expressed uncertainty even when her answer was correct. Task 8 also led to a similar pattern of manual computation. B1 said aloud “I guess I could add them. […] Is there a way to select all, or is that just it?” Similar sentiments were expressed by B4 and B5. B4 even wished for a similar feature as in Excel where points are added automatically: “just like what a[n] Excel table will do… so that I don’t have to calculate by myself.” This highlighted how unused or unfamiliar interaction techniques in Power BI led participants to revert to manual addition. Overall, participants did not use keyboard-and-mouse interactions, including selecting data points to obtain aggregated values.
6.3.3. Hard Tasks
Similar to medium tasks, hard tasks were also computationally demanding, and participants were also uncertain about their answers. Participants frequently questioned their answers (B3, B5, and B6) and, at times, provided incorrect responses or were unable to complete the tasks. Each participant approached Task 3 differently. B1 discovered and used the drill-down interaction on the map visual to solve the task, whereas B2, B3, and B4 relied on manual calculations rather than using the available interactivity. Task 5, which also required drill-down interaction, was solved by only half of the participants. B3 and B4 located the drill-down feature but were unable to use it effectively to reach the correct answer. B5 was unable to solve the task altogether; she waited for additional explanations on hover and appeared to expect the system to surface more guidance automatically. B6 often consulted the user guide to find answers and to see which visualizations were explained during the task-solving phase.
Figure 6. Comparing the results between Diana and the baseline condition. Overall results for each task for participants from Diana condition (D1-D6) and from the baseline condition (B1 - B6). Results indicate that the participants from the baseline condition performed worse than the participants from the Diana condition.
This table compares task outcomes for participants D1–D6 and B1–B6 across ten tasks, with icons indicating whether each task was solved correctly, solved incorrectly, or not solved. The pattern across the table shows that Hard tasks tend to yield the fewest correct solutions and the highest rates of failure or incomplete attempts. Medium tasks produce a mixture of correct and incorrect results, while Easy tasks are generally completed successfully by both groups. Although the difficulty gradient is consistent, the two participant groups differ in the types of tasks they handle well, and Diana participants perform better on interpretive tasks, while Baseline participants appear more successful on interaction-focused tasks.
6.4. Comparison Between Diana and Static Dashboard Guide
To better understand the effect of the onboarding method, we compared task performance (as shown in Figure 6), answer correctness, and participants’ overall experience using Diana with those who relied solely on the static dashboard guide.
6.4.1. Task Accuracy and Completeness
Participants who used Diana performed better overall than those relying solely on the user guide. As task difficulty and cardinality increased, baseline condition participants struggled both to interpret the dashboard and to arrive at correct answers. We distinguish between tasks participants could not solve or skipped (completeness) and those answered incorrectly (accuracy). In both study groups, at least one task could not be solved by a participant. In the Diana group, this occurred only once: D1 was unable to solve Task 3. In contrast, in the baseline condition group, five tasks were left unsolved: one each by B2, B3, and B6, and two by B1. For example, B6 could not complete Task 8 because she did not realize that, in her earlier interactions, she had applied filters, which had caused the map to hide relevant data points.
Incorrect answers were also more prevalent in the static dashboard guide group. Across participants B1 –B6, a total of ten tasks were answered incorrectly: five by B5, two each by B4 and B6, and one by B2. Despite differences in visualization and dashboard expertise, every participant in the guide group had at least one task that they either could not solve or answered incorrectly.
6.4.2. Feeling of Uncertainty
Although confidence was not an explicit study measure, we observed a notable contrast between the two groups in how they expressed uncertainty while solving tasks. Participants using Diana occasionally sought confirmation of their answers (e.g., D2 and D3) and were successful in receiving it via one or more modalities.
In contrast, participants relying on the baseline condition expressed persistent uncertainty throughout the study. They frequently used phrases such as “not sure” or “not right, but…” while reasoning aloud. For instance, during Task 6, despite eventually providing the correct answer, B6 remarked, “Okay, it’s definitely not right. I don’t think so. But if this is what I have here…” Similar hesitations were observed for B2, B3, B4, B5, and B6, with each of them voicing uncertainty at least once during the session.
6.4.3. Underexplored Interactivity and Reliance on Manual Calculations
Participants using Diana learned a range of interaction techniques through onboarding support and demonstrated a learning effect by Task 9. In contrast, participants in the baseline group struggled with basic and intermediate interactivity throughout the study. All participants in the baseline condition resorted to manually adding data-point values often performing arithmetic on paper because they did not understand or use the available keyboard-based interactions. Microsoft Power BI dashboards require a combination of mouse and keyboard interactions (e.g., multi-selecting values, drilling down, removing filters), yet these participants relied almost entirely on mouse hovering or visual inspection.
6.5. Reflections on Diana
@@
In the post-task interview, participants (D1-D6) reflected on three main themes: their choice of modality, their experience with the onboarding, and the comparison to previous onboarding experiences.
6.5.1. Preferred Modalities
As participants conducted the user study using Diana, their preferred modality became clear. Despite varying levels of prior experience, all participants experimented with speech interaction. Even those with no prior exposure (D1, D2, D6) adopted it during the study, with D2 initially hesitant but later describing it as “comfortable, almost not robotish.” Voice was valued for its speed and natural phrasing, while chat was preferred for step-by-step guidance (D1 and D5). Mouse interactions, such as lasso selection, were consistently used to define a focus before turning to another modality (D3 and D4).
@@ 6.5.1. Preferred Modalities
Participants also found multiple modalities useful, as they allowed them to adapt depending on time, task, or personal style. D1 summarized this flexibility: “If you are in a hurry, you can just talk to it, and it goes faster. If you have time to sit down and think, then you can use the chat and the radial menu… But I would probably use writing at most, because I’m old-fashioned.”
Across modalities, the visual highlight feature was universally appreciated to orient user attention.
6.5.2. Strengths of Voice-based Interaction and Visual Highlights Feature
@@
Speech interaction was well received, even among those without prior experience. D2 remarked that “it was very comfortable voice, I could hear that it was mechanical, but it was not that mechanical.” He also appreciated the real-time response, and when paired with the visual highlighting feedback, it helped him see what Diana was saying in response to his question (DG4 (orientation)). As D2 was an expert, he could verify the correctness of the answer provided by Diana. As the answer was phrased naturally and accurately, he said “and when it did that [gave him the right steps to the answer] it had my trust, that I could actually ask the voice interaction a question.” D3 described the voice modality and the response from Diana as informative and that he “really [had] fun interacting with her,” while D4 appreciated that speech “doesn’t stop when you are doing something else on the dashboard.” The mouse modality, in the form of lasso selection, was also valued when combined with voice, as the combination helped participants anchor their queries to specific visuals.
6.5.3. Underused Modalities and Features
The radial menu was the least used feature. While D2 found it helpful for confirmation, most others described it as unintuitive, overly detailed, or unfamiliar: “I didn’t know what specifically to use it for […] the surface things were good enough” (D4). The user guide was largely ignored, with D1 and D3 admitting they had forgotten it existed. D6 said she would have preferred to read it if time allowed. This indicates that static documentation can be easily overshadowed by interactive support. Although useful for step-by-step guidance, the chat via keyboard was sometimes described as slow or effortful (“chatting took too long to write compared to just speaking,” D2).
6.5.4. Comparison to Prior Onboarding.
@@ 6.5.3. Underused Modalities and Features
Compared to step-by-step onboarding methods, Diana was described as more flexible and empowering. D4 compared it to video game tutorials where you are forced to proceed linearly: “You have to explore one step at a time… Here you have the freedom to onboard yourself the way you want to.” D1 echoed this, preferring to ask for help when needed rather than “click all over” through a fixed tutorial (DG5 (guidance & autonomy)). For D2, who regularly onboards colleagues through videos and one-on-one sessions, Diana represented a potential time-saver. Although he values personal interaction, he saw the system as a “virtual project finance controller” (DG3 (human-like)) that could reduce his workload.
7. Discussion
@@
7.1. Speech Interaction as the Human Onboarder Assistant
In the main study, participants frequently preferred speech over mouse or keyboard-based interactions for onboarding help and follow-up questions. This result matches the findings of León et al. (Molina León et al., 2024) on modality preferences for exploration tasks. This is also in line with the findings from Zhan et al. (Zhan et al., 2024) who studied trust in healthcare voice assistants and suggested that perceived usefulness of a system is a strong indicator of the trust a user has in the system.
Rather than replacing human onboarders, Diana assists and augments their capabilities by handling routine questions such as how do I…? that are time-consuming to document and maintain. In the context of onboarding, the result also aligns with one of our initial goals of providing human-like support (DG3 (human-like)). The findings suggest that participants appreciated real-time responses and being less dependent on asking a human expert for help.
In our formative study, we also came across a surprising result: the Realtime API (OpenAI, 2023b) detected that the participant was not a native English speaker and switched to another language that the person understood. While unexpected, the person was positively surprised by the detection. Furthermore, other participants wished for natural language interaction in their native language. While this is an option with current large language models, it may go beyond the capabilities of many human onboarders.
7.2. Multimodal Interaction for Onboarding
@@
The less use of the contextual menu could also suggest that sophisticated visual encodings for help are not necessary for onboarding, as they were perceived as advanced or quite detailed. But, they would be extremely helpful at later stages after users have been onboarded, to allow them to access more fine-grained information, as echoed by D1.
7.3. Limited Adoption of Radial Menu
Even though the radial menu was designed to provide structured and faster guidance compared to other LLM-powered onboarding options, such as chat and speech interaction, it was the least used feature overall. Participants reported several reasons for this, such as unintuitive (D3), too much information (D2, D4), and generally time-consuming to find answers (D1). In addition, the unfamiliar design (D4) and the presence of technical terms may have hindered adoption. Unlike the LLM-powered onboarding features, which provided help tailored to users’ needs, the radial menu required participants to actively explore and discover relevant information, which could also explain its low usage. We plan to improve the radial menu design based on feedback from the user study.
7.4. Using Large Language Models for Onboarding
@@
While the natural interaction via voice and text was successful, it required a substantial effort at the technical level to minimize the number of hallucinations. The information sent to the LLMs was carefully designed based on available REST API information and heuristics to ensure that information from the REST APIs is not wildly extrapolated. The prompts were fine-tuned several times throughout the formative study. We were able to significantly reduce hallucinations in both chat- and voice-based interactions, but participants still encountered issues when they highlighted a specific section of the dashboard and asked for guidance via natural language.
@@
Additionally, because Diana has no access to the data, it can be used in any domain with ease, as it only provides information about the high-level aspects of the visualization. This was an intentional design choice motivated by privacy and security concerns, and it also prevented the LLM from giving direct answers, which participants attempted to obtain on several occasions. It can infer trends and drivers, however, it cannot give detailed data-level information that is usually critical in a business context. With controlled access to data values and their explanations, the system could provide richer, more contextualized guidance during dashboard exploration.
7.5. Limitations
We examine the limitations of Diana with respect to its design, generalizability, deployment, and practical considerations, including cost. We also outline the limitations of our user study.
7.5.1. Design Limitations
The design limitations of Diana became evident during the user study. The low usage of the radial menu calls for a redesign, including simplifying its hierarchy or reducing the amount of information shown at once. Beyond simplification, different forms of contextual or on-visual menus may also need to be explored. Additionally, sometimes the visual highlights were not synchronized with the voice output, causing confusion.
7.5.2. Generalizability
Diana supports Microsoft Power BI dashboards with default visuals. Many of its core design concepts—such as the Vue-based frontend components, multimodal interaction features, and an LLM-backed backend with structured prompting—could be generalized to other dashboard platforms and custom visuals. The design concepts may generalize for custom visuals and new dashboard platforms, however technical implementation would require some engineering effort, such as:
We developed Diana specifically for Microsoft Power BI, and it currently works with any dashboard with default visuals. It can also support custom visuals, but only if developers supply the necessary metadata, as Power BI does not expose this information by default.
The frontend is implemented as a set of Vue-based components. Therefore, many of the multimodal interaction features could be reused in other dashboarding environments. The backend has the LLM integration and the prompting structure, which are also conceptually transferable.
While the design concepts may generalize, the technical implementation would require some adaptation to function in other dashboard systems.
Cross-Platform Deployment
Integrating Diana to non-Power BI dashboards may require some engineering effort, such as:
  • Metadata extraction: For new platforms, visual metadata would either need to be structured according to the current implementation or the implementation adapted to the new platform’s metadata format.
  • Interaction support: Power BI uses keyboard and mouse interactions, which may differ or be absent on other platforms. Therefore, prompts and interaction handling would need to be adjusted accordingly.
Cross-platform deployment could enable broader applicability of Diana, but we have not yet empirically validated it beyond Power BI. Future work could explore prototyping on alternative dashboard platforms to confirm feasibility and usability.
7.5.3. Cost Considerations
We also relied on OpenAI APIs, including the Realtime API, which incurs usage-based costs. The Realtime API (OpenAI, 2023b) can become expensive depending on interaction volume. As AI models continue to improve, these costs may decrease, and more affordable alternatives may emerge. However, the current API pricing remains a practical concern for the widespread deployment and generalizability of Diana.
7.5.4. User Study Limitations
The user study was conducted on a single dashboard to avoid bias, ensuring consistent difficulty and controlled cardinality (i.e., the number of views a task required). However, this choice limits the generalizability of the findings to that specific dashboard. Additionally, the two study groups differed slightly in their prior expertise, which may have influenced performance differences. Specifically, participants in the Diana condition reported slightly greater prior experience with dashboards compared to those in the static dashboard guide group.
Future work could investigate learning improvement and time efficiency more systematically. This would require a longitudinal study in which Diana is deployed over an extended period and compared against multiple baseline onboarding methods, such as static guides, video tutorials, or step-by-step tutorials, to provide quantitative evidence of its effectiveness.
8. Conclusion
@@
We designed and developed Diana, a multimodal dashboard assistant that combines voice, text, in-situ visual highlights, and contextual help to support users during dashboard onboarding. In an exploratory user study, we found that participants using Diana achieved higher task accuracy and completeness than those using the static dashboard guide. Participants in the Diana condition used all interaction modalities and preferred voice-based interaction over keyboard- and mouse-based interaction. Our findings suggest that LLMs are useful for onboarding tasks when they are paired with visual affordances. It also encouraged users to ask for help without compromising their autonomy. We suggest that onboarding systems should be multimodal by design and offer a low-friction path for users to onboard themselves on dashboards.
Acknowledgements.